?On December 31, 2015, Bonita Co. performed environmental consulting services for Hayduke Co. Hayduke was short of cash, and Bonita Co. agreed to accept a $346,500 zero-interest-bearing note due December 31, 2017, as payment in full. Hayduke is somewhat of a credit risk and typically borrows funds at a rate of 11%. Bonita is much more creditworthy and has various lines of credit at 6%.

?On December 31, 2015, Bonita Co. performed enviro
Answer #1

The answer has been presented in the supporting sheet. All the parts ha sbeen solved with detailed explanation and format.

2 Answer Firstl we have to calculate the present value of bond = Face value / (1+i)^n =346500/(1.11)^2 =$ 281227 3 4 Credit D